OpenStudy (anonymous):

given the center and the radius of a circle, write the equation describing the circle. (0,0),r=8 (x-0)^2+(y-4)^2=81 x^2+(y-4)^2=81

OpenStudy (anonymous):

(0,0) r = 8 would be x^2 + y^2 = 64 because (x - 0)^2 + (y - 0)^2 = 8^2

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

I don't really know what you are asking.

OpenStudy (anonymous):

whats the answer and how did you find it sweety?

OpenStudy (anonymous):

What are the last two equations there for? The questions says write the equation of a circle with center (0,0) and radius 8... The answer to that is x^2 + y^2 = 64

OpenStudy (anonymous):

i dont know thats what it says